# Tensor network methods for quantum-inspired image processing and classical optics

### v1

@@ V Wave optics and quantum mechanics
Interestingly, many other optical propagators might be good candidates for finite-rank MPO approximations, as they often respect strict conditions of smoothness, locality and symmetry. Previously we have considered the far-field regime described by the Fraunhofer propagation (the Fourier transform), a similar analysis can be generalized to the near-field regime, namely the Fresnel propagation which is valid for small distance of propagation ($z\ll D$) in the paraxial approximation, and more generally to the free propagation at any distance.
V.1 Free optical propagation
@@
Figure 15: Tensor representation of the optical propagators
As a warm-up exercise, before tackling the more difficult task of systematically constructing low-rank matrix product operators for optical propagators, we first employ a brute-force method (computationally infeasible for large $n$) to diagnose the structure of these operators. In Fig. 16, we show how to construct MPO’s by SVD in a similar fashion we used for MPS previously.
@@
The propagator Fig. 15 described as a MPO $\hat{U}_{AS}$, can not be written as separable operator $U^{(x)}\otimes U^{(y)}$ because of the square root in Eq. (15) causing couplings between the frequencies. The amount of coupling (entanglement) can nevertheless be quantify by organizing the transfer function into a matrix product operator with a fused representation and looking at the rank between the two halves of the operator (see Fig. 17).
Figure 17: Bond dimensions along a chain of 12 qubits for the $2d$ Fresnel and $2d$ ASTF matrix product operators in a fused representation. In this setting, the left half of the chain corresponds to the $x$-component and the right half corresponds to the $y$-component.
@@
V.2 Fresnel propagator and the free quantum particle
@@
The addition of a potential $V(x)$ in the former would imply a heterogenous refraction index $n(x)$ in the later. As said earlier, the Fresnel/free particle propagator has, like the Fourier transform, a separable tensor product structure in space, so we can focus on its $1d$ discretization in the following.
V.2.1 Discretization of the Hamiltonian
@@
V.2.2 Hamiltonian in the qubit basis
Having wrote the Fresnel transform as the unitary propagation of a quantum Hamiltonian of $L$ hopping particles on a lattice, we can proceed to compress the Hilbert space acting on $n=\log{L}$ qubits. Consider an $n$-qubit computational basis labeled by bitstrings $\mathbf{s}=(s_{1},s_{2},\ldots,s_{n})$ where each $s_{j}\in\{0,1\}$. Define the number $x_{\mathbf{s}}=\sum_{j=1}^{n}s_{j}2^{-j}\in[0,1)$, we then need to express the dispersion relation of the diagonalized Hamiltonian such that $\hat{H}$ remains diagonal in the computational basis. Any diagonal hamiltonian (or operator for that matter) can be expanded in Pauli-strings of any size
@@ V.2.2 Hamiltonian in the qubit basis
This decomposition is the Walsh–Hadamard transform (which is a generalized Fourier transform over the abelian group $\mathbb{Z}_{2}^{n}$), it can be computed very efficiently for large-system with the fast Walsh–Hadamard transform (see for example Georges et al. (2025) and references within). Generally speaking, diagonal operators are crucially important in quantum computing as there are key elements of many quantum algorithms Bullock and Markov (2003, 2008); Welch et al. (2014); Nakata and Murao (2014).
As a brief illustration of the method for $n=2$. The possible Pauli-$Z$ tensor products are
$\{\mathbb{I}\otimes\mathbb{I},\ Z\otimes\mathbb{I},\ \mathbb{I}\otimes Z,\ Z\otimes Z\}.$ (32)
[…]
$c_{a}=\frac{1}{4}\sum_{x=0}^{3}d_{x}\chi_{a}(x),\quad a\in\{00,01,10,11\},$ (33)
with $\chi_{a}(x)=(-1)^{a.x}$ for the group $\mathbb{Z}_{2}^{2}$. This provides the coefficients for the decomposition $\hat{D}=\sum_{a\in\{0,1\}^{2}}c_{a}Z^{a}$.
In our specific case $\epsilon(\mathbf{s})=\cos(2\pi x_{\mathbf{s}})$, a standard calculation (cf. appendix) gives the coefficients $c_{S}\in\mathbb{R}$ equal to
@@
and $e_{0}=1$  [^8] Actually, all the moments can be written using those polynomials, giving access to $U(t)$ as an exact moment expansion.. Thanks to the convergence properties of $c_{S}$, we can show (more details in the appendix) that the asymptotic behavior with respect to $P$ is
@@
This decay is super-exponential in $P$ and it shows that the amplitudes of Pauli strings do not contribute at all past a certain string length. We plot in Fig. 19 this asymptotic behavior and compared it to the exact formula computed with the fast Walsch-Hadamard transform.
@@
V.2.3 Unitary evolution, quantum circuit and MPO
@@
Similarly to the Fourier transform MPO, we have successfully showed that the Fresnel transform in Fourier space also admits a low rank structure in Fourier space. The real-space operator is also low-rank since it is a product of low-rank MPO (see Fig. 18). The numerical method developed here is very generic and is expected to work with any diagonal operator. Of course the maximal dimension of the MPO will grow with the complexity of the function $\mathbf{\epsilon}(\mathbf{s})$, but can always be controlled by truncating the terms in the Pauli expansion. As long as the behavior of $\overline{c}_{p}$ is sufficiently decreasing, the quantum circuit and the MPO associated should stay numerically tractable for large value of $n_{\mathrm{qubit}}$. Additionally, more subtle and efficient ways, such as recursive decomposition into multiplexed $R_{z}$ gates Shende et al. (2005); Bullock and Markov (2003) allows to construct the MPO using $\mathcal{O}(n)$ gates. In the next section, we will introduce another method that is more mathematically explicit and offers some advantages in particular circumstances.
V.2.4 Closed-expression for the diagonal MPO using Bessel expansion
@@
with $\theta_{j}=\frac{2\pi}{2^{j}}$ and where $J_{p}(t)$ is the $p$-th Bessel function of the first kind. The last formula is directly a matrix product operator as a superposition of plane waves weighted by Bessel-function coefficients. In practice, the infinite sum over $p$ is truncated to $|p|\leq p_{\rm max}$. Let $\chi=2p_{\rm max}+1$ be the bond dimension, corresponding to the $p$ index. Define local tensors $W_{j}$ with physical index $s_{j}$ and bond indices $p_{\rm in},p_{\rm out}$ as
@@
V.3 Angular spectrum propagator as a free relativistic quantum particle
@@
This yields the discrete Dirac Hamiltonian [^10] The infamous fermion doubling problem will not pose a significant problem in our context
@@
for $\quad\mu\geq 1$ with $\gamma^{-1}=\sqrt{1-\beta^{2}}$, and $(a)_{\mu}=a(a+1)(a+2)\cdots(a+\mu-1)$ and where $\lambda=\beta/(1+\gamma^{-1})$. For fixed $\mu$, the integrals converge absolutely. As $\mu\to\infty$, $a_{\mu}$ decays exponentially because the function $\sqrt{1-\beta^{2}\sin^{2}\phi}$ is analytic in a strip around the real axis since $0<\beta<1$. Because of the exponentially decreasing behavior of the Pauli string coefficients, the MPO also admits a very low-rank representation. We do not show the quantum unitary circuit or the MPO implementation, as they closely mirror the Fresnel propagator. In this study, the dispersion relations are analytic functions of $k$ implying exponentially fast decrease of their Hadamard coefficients. Mathematically it sounds reasonable to assert that any Hamiltonian with an analytic/holomorphic dispersion relation should admit a low-rank MPO representation. It would be very important to precisely quantify this behavior more rigorously for a larger class of functions. Alternatively, a similar Bessel-type expansion can also be found in this case, leading to a closed expression for this MPO, although we do not show details here as they do not add value to the discussion.
Additionally, it would be informative to compare the rank of the resulting MPO with that obtained from a more TCI-based implementation. Also, another quantum-inspired method, which does not rely on the Fourier transform, has been recently developed Gidi et al. (2025) to solve quantum free evolution. Such comparisons, along with more realistic applications, will be the subject of future research. Surprisingly enough, although we have dealt with two operators that are diagonalizable by Fourier transform, we did not construct the Fourier transform MPO by this method as its associated Hamiltonian $\hat{\mathcal{H}}=\frac{1}{2}(\hat{p}^{2}+\hat{x}^{2})$ is not diagonalizable by Fourier transform (as $\hat{H}$ is not translational invariant), but rather by the basis of discrete Hermite functions. It would be very interesting to calculate an MPO representation using this Hamiltonian formulation and compare with Chen et al. (2023).
VI Conclusions and perspectives
@@
We showed how Fourier optics can be cast as a form of quantum mechanics, with qubits whose evolution is dictated by unitary operators that can be represented as matrix product operators, allowing efficient computations. In the future, we plan to extend this work to broader optical applications and explore algorithms for optical inverse problems—such as phase retrieval, denoising, and deconvolution—which have numerous applications in astronomy, earth observation, and microscopy.
@@
Acknowledgements I am deeply grateful to Bernhard Jobst for discussions in the early stage of this paper, and to all the contributors to the ITensor library ecosystem Fishman et al. (2022). All simulations have been carried out on a 2015 macbook pro 16GB.
Appendix A Pauli string exact formula

### v2

@@ VI Wave optics and quantum mechanics
Interestingly, many other optical propagators might be good candidates for finite-rank MPO approximations, as they often respect strict conditions of smoothness, locality and symmetry. Previously we have considered the far-field regime described by the Fraunhofer propagation (the Fourier transform), a similar analysis can be generalized to the near-field regime, namely the Fresnel propagation which is valid for small distance of propagation in the paraxial approximation, and more generally to the free propagation at any distance.
VI.1 Free optical propagation
@@
Figure 18: Spatial separability of the optical propagators in tensor notation.
As a warm-up exercise, before tackling the more difficult task of systematically constructing low-rank matrix product operators for optical propagators, we first employ a brute-force method (computationally infeasible for large $n$) to diagnose the structure of these operators. In Fig. 19, we show how to construct MPO by SVD in a similar fashion we used for MPS previously.
@@
The propagator Fig. 18 described as a MPO $\hat{U}_{AS}$, can not be written as separable operator $U^{(x)}\otimes U^{(y)}$ due to the the square root in Eq. (16) causing couplings between the frequencies. The amount of coupling (entanglement) can nevertheless be quantified by organizing the transfer function into a matrix product operator with a fused representation and looking at the rank between the two halves of the operator (see Fig. 20).
Figure 20: Bond dimensions $\xi$ along a chain of 12 qubits for the $2d$ Fresnel and $2d$ ASTF matrix product operators in a fused representation for $z=0.1m$. In this setting, the left half of the chain corresponds to the $x$-component and the right half corresponds to the $y$-component.
@@
VI.2 Fresnel propagator and the free quantum particle
@@
The addition of a potential $V(x)$ in the former would imply a heterogenous refraction index $n(x)$ in the later. We shall use the time representation (rather than $z$) in the following sections to better match standard quantum hamiltonian evolution formalism. As said earlier, the Fresnel/free particle propagator has, like the Fourier transform, a separable tensor product structure in space, so we can focus on its $1d$ discretization in the following.
VI.2.1 Discretization of the Hamiltonian
@@
VI.2.2 Hamiltonian in the qubit basis
Having written the Fresnel transform as the unitary propagation of a quantum Hamiltonian of $L$ hopping particles on a lattice, we can proceed to compress the Hilbert space acting on $n=\log_{2}{L}$ qubits. Consider an $n$-qubit computational basis labeled by bitstrings $\mathbf{s}=(s_{1},s_{2},\ldots,s_{n})$ where each $s_{j}\in\{0,1\}$. Define the number $x_{\mathbf{s}}=\sum_{j=1}^{n}s_{j}2^{-j}\in[0,1)$, we then need to express the dispersion relation of the diagonalized Hamiltonian such that $\hat{H}$ remains diagonal in the computational basis. Any diagonal hamiltonian (or operator for that matter) can be expanded in Pauli-strings of any size
@@ VI.2.2 Hamiltonian in the qubit basis
This decomposition is the Walsh–Hadamard transform (which is a generalized Fourier transform over the abelian group $\mathbb{Z}_{2}^{n}$), it can be computed very efficiently for large-system with the fast Walsh–Hadamard transform (see for example Georges et al. (2025) and references within). Generally speaking, diagonal operators are crucially important in quantum computing as they are key elements of many quantum algorithms Bullock and Markov (2003, 2008); Welch et al. (2014); Nakata and Murao (2014).
In our specific case $\epsilon(\mathbf{s})=\cos(2\pi x_{\mathbf{s}})$, a standard calculation (cf. appendix B) gives the coefficients $c_{S}\in\mathbb{R}$ equal to
@@
and $e_{0}=1$  [^17] Actually, all the moments can be written using those polynomials, giving access to $U(t)$ as an exact moment expansion.. Thanks to the convergence properties of $c_{S}$, we can show (more details in the appendix B) that the asymptotic behavior with respect to $P$ is
@@
This decay is super-exponential in $P$ and it shows that the amplitudes of Pauli strings do not contribute at all past a certain string length. We plot in Fig. 22 this asymptotic behavior and compare it to the exact formula computed with the fast Walsch-Hadamard transform.
@@
VI.2.3 Unitary evolution, quantum circuit and MPO
@@
Similarly to the Fourier transform MPO, we have successfully showed that the Fresnel transform in Fourier space also admits a low rank structure in Fourier space. The real-space operator is also low-rank since it is a product of low-rank MPO (see Fig. 21). The numerical method developed here is very generic and is expected to work with any diagonal operator. Of course the maximal rank of the MPO will grow with the complexity of the function $\mathbf{\epsilon}(\mathbf{s})$, but can always be controlled by truncating the terms in the Pauli expansion. As long as the behavior of $\overline{c}_{p}$ is sufficiently decreasing, the quantum circuit and the MPO associated should stay numerically tractable for large value of $n_{\mathrm{qubit}}$. Additionally, more subtle and efficient ways, such as recursive decomposition into multiplexed $R_{z}$ gates Shende et al. (2005); Bullock and Markov (2003) allows to construct the MPO using $\mathcal{O}(n)$ gates. In the next section, we will introduce another method that is more mathematically explicit and offers some advantages in particular circumstances.
VI.2.4 Closed-expression for the diagonal MPO using Bessel expansion
@@
with $\theta_{j}=\frac{2\pi}{2^{j}}$ and where $J_{p}(t)$ is the $p$-th Bessel function of the first kind. The last formula is directly a matrix product operator as a superposition of plane waves weighted by Bessel coefficients. In practice, the infinite sum over $p$ is truncated to $|p|\leq p_{\rm max}$. Let $\chi=2p_{\rm max}+1$ be the bond dimension, corresponding to the $p$ index. Define local tensors $W_{j}$ with physical index $s_{j}$ and bond indices $p_{\rm in},p_{\rm out}$ as
@@
VI.3 Angular spectrum propagator as a free relativistic quantum particle
@@
This yields the discrete Dirac Hamiltonian [^19] The infamous fermion doubling problem will not pose a significant problem in our context, since we can choose one branch arbitrarily.
@@
for $\quad\mu\geq 1$ with $\gamma^{-1}=\sqrt{1-\beta^{2}}$, and $(a)_{\mu}=a(a+1)(a+2)\cdots(a+\mu-1)$ and where $\lambda=\beta/(1+\gamma^{-1})$. For fixed $\mu$, the integrals converge absolutely. As $\mu\to\infty$, $a_{\mu}$ decays exponentially because the function $\sqrt{1-\beta^{2}\sin^{2}\phi}$ is analytic in a strip around the real axis since $0<\beta<1$. Because of the exponentially decreasing behavior of the Pauli string coefficients, the MPO also admits a very low-rank representation. We do not show the quantum unitary circuit nor the MPO implementation, as they closely mirror the Fresnel propagator.
In this study, the dispersion relations are analytic functions of $k$ implying exponentially fast decrease of their Hadamard coefficients. Mathematically it sounds reasonable to assert that any Hamiltonian with an analytic/holomorphic dispersion relation should admit a low-rank MPO representation. It would be very important to precisely quantify this behavior more rigorously for a larger class of functions. Alternatively, a similar Bessel-type expansion can also be found in this case, leading to a closed expression for this MPO, although we do not show details here as they do not add value to the discussion.
Additionally, it would be informative to compare the rank of the resulting MPO with that obtained from a more TCI-based implementation. Also, another quantum-inspired method, which does not rely on the Fourier transform, has been recently developed Gidi et al. (2025) to solve quantum free evolution. Such comparisons, along with more realistic applications, will be the subject of future research. Moreover, many other transforms that are diagonalizable in Fourier could be study within this framework. One of the most famous transform that is especially present in image processing, optics and quantum physics is the periodic Hilbert transform. In Fourier, the Hilbert operator is a simple chiral phase operator $\pm i$ which can be written $\hat{U}_{\mathrm{QFT}}^{\dagger}\hat{D}\hat{U}_{\mathrm{QFT}}$, with $\hat{D}=-i\hat{Z}_{n}$ being a rank-1 MPO. A more general analysis of Fourier operators could open new possibilities at the interface of quantum computing, filter theory and tensor networks.
Surprisingly enough, although we have dealt with two operators that are diagonalizable by Fourier transform, we did not construct the Fourier transform MPO by this method as its associated Hamiltonian $\hat{\mathcal{H}}=\frac{1}{2}(\hat{p}^{2}+\hat{x}^{2})$ is not diagonalizable by Fourier transform (as $\hat{\mathcal{H}}$ is not translational invariant), but rather by the basis of discrete Hermite functions. It would be very interesting to calculate an MPO representation using this Hamiltonian formulation and compare with Chen et al. (2023).
VII Conclusions and perspectives
@@
We showed how Fourier optics can be mapped to quantum mechanics, with qubits whose evolution is dictated by unitary operators that can be represented as matrix product operators, allowing efficient computations. In the future, we plan to extend this work to broader optical applications and explore algorithms for optical inverse problems—such as phase retrieval, denoising, and deconvolution—which have numerous applications in astronomy, earth observation, and microscopy.
@@
Acknowledgements I am deeply grateful to Bernhard Jobst for discussions in the early stage of this paper, to all the contributors to the ITensor library ecosystem Fishman et al. (2022) and to one of the referee for its careful reading and constructive criticism. All simulations have been carried out on a 2015 macbook pro 16GB. The JPG picture used in this article will be made publicly available or can be obtained from the author upon request.
Appendix A Matrix product states and binary tensor trees
In this appendix, we show explicitly how to construct MPS and tensor trees (figure below) that we used in the main text.
Figure 28: Example of a matrix product state with $n=8$ and bulk binary tensor tree with $n=14$. The construction is shown below
In both cases, the construction is based on successive singular value decompositions along the link of a network. The full $n$-qubit state vector $|\psi\rangle\in\mathcal{H}$, whether we choose the FRQI or quantics encodings is
$|\psi\rangle=\sum_{s_{1},s_{2},...,s_{n}\in\mathbb{Z}_{2}}\psi_{s_{1}s_{2}...s_{n}}|s_{1}s_{2}...s_{n}\rangle,$ (65)
where the tensor $\psi_{s_{1}s_{2}\dots s_{n}}$ is represented by the purple $n$-leg object in Fig. 7.
Matrix product states
The matrix decomposition outlined below is a generalization of the singular value decomposition and can be used to decompose a generic tensor into constituent tensors of lower order. Every bond corresponds to a Schmidt decomposition of the Hilbert space $\mathcal{H}=\mathcal{H}_{[1,2,...,k]}\otimes\mathcal{H}_{[k+1,k+2,...,n]}$. The Schmidt decomposition across the cut can be written
$|\psi\rangle=\sum_{\alpha=1}^{\chi_{k}}\lambda^{[k]}_{\alpha}|\alpha\rangle_{[1,2,..,k]}\otimes|\alpha\rangle_{[k+1,k+2,..,n]}$ (66)
with orthonormality of the Schmidt bases generated by contracting left ($|\alpha^{\prime}\rangle_{[1,2,..,k]}$) and right-canonical tensors ($|\alpha^{\prime}\rangle_{[1,2,..,k]}$) such that $\langle\alpha|\alpha^{\prime}\rangle_{[1,2,..,k]}=\delta_{\alpha\alpha^{\prime}}$ and $\langle\alpha|\alpha^{\prime}\rangle_{[k+1,k+2,..,n]}=\delta_{\alpha\alpha^{\prime}}$. Expressing the Schmidt basis states in the computational basis
$|\alpha\rangle_{[1,...,k]}$ $=\sum_{s_{1},...,s_{k}}L^{[1..k]}_{s_{1}...s_{k},\alpha}|s_{1}...s_{k}\rangle,$ (67)
$|\alpha\rangle_{[k+1,...,n]}$ $=\sum_{s_{k+1},...,s_{n}}R^{[k+1..n]}_{\alpha,s_{k+1}...s_{n}}|s_{k+1}...s_{n}\rangle.$ (68)
Substituting back into the Schmidt decomposition gives
$|\psi\rangle=\sum_{s_{1},\dots,s_{n}}\sum_{\alpha=1}^{\chi_{k}}\lambda^{[k]}_{\alpha}L^{[1..k]}_{s_{1}...s_{k},\alpha}R^{[k+1..n]}_{\alpha,s_{k+1}...s_{n}}|s_{1}...s_{n}\rangle.$ (69)
Factorizing left and right Schmidt states recursively
$|\alpha\rangle_{[1,...,k]}=\sum_{\begin{subarray}{c}\alpha_{1},\dots,\alpha_{k-1}\\ s_{1},...,s_{k}\end{subarray}}A^{[1]}_{s_{1},\alpha_{1}}A^{[2]}_{\alpha_{1},s_{2},\alpha_{2}}..A^{[k]}_{\alpha_{k-1},s_{k},\alpha}|s_{1}...s_{k}\rangle,$
$|\alpha\rangle_{[k+1,\dots,n]}=\sum_{\begin{subarray}{c}\alpha_{k+1},...,\alpha_{n-1}\\ s_{k+1},...,s_{n}\end{subarray}}A^{[k+1]}_{\alpha,s_{k+1},\alpha_{k+1}}..A^{[n]}_{\alpha_{n-1},s_{n}}|s_{k+1}...s_{n}\rangle,$
leads to the usual form of the factorization of components
$\psi_{s_{1}s_{2}\dots s_{n}}=\sum_{\alpha_{1},\dots,\alpha_{n-1}}A^{[1]}_{s_{1},\alpha_{1}}A^{[2]}_{\alpha_{1},s_{2},\alpha_{2}}..A^{[n]}_{\alpha_{n-1},s_{n}}$, (70)
which is represented diagrammatically by the blue object in Fig. 7. Depending on the scaling of $\chi_{k}$ with $n$, the MPS can efficiently represent $\psi_{s_{1}s_{2}\dots s_{n}}$.
Binary tensor tree with physical sites in the bulk
Once fixed a binary tree topology, every SVD corresponds to a tree cut which can separate non-contiguous qubits in the original ordering $\{1,2,...,k\}|\{k+1,k+2,...,n\}$. We can decompose the full tensor into a binary tree of lower-order tensors using a sequence of Schmidt decompositions corresponding to the cuts defined by the edges of the tree. Every edge $e$ of the tree defines a bipartition of the Hilbert space
$\mathcal{H}=\mathcal{H}_{\text{subtree}(e)}\otimes\mathcal{H}_{\text{rest}(e)}.$ (71)
Let us explicite the construction for the bulk tree (see Fig. 10). The Schmidt decomposition across the edge $e$ reads
$|\psi\rangle=\sum_{\alpha=1}^{\chi_{e}}\lambda^{(e)}_{\alpha}|\alpha\rangle_{\text{subtree}(e)}\otimes|\alpha\rangle_{\text{rest}(e)},$ (72)
with orthonormality conditions
$\langle\alpha|\alpha^{\prime}\rangle_{\text{subtree}(e)}=\delta_{\alpha\alpha^{\prime}},\qquad\langle\alpha|\alpha^{\prime}\rangle_{\text{rest}(e)}=\delta_{\alpha\alpha^{\prime}}.$ (73)
Expressing the Schmidt basis states in the computational basis, we have
$|\alpha\rangle_{\text{subtree}(e)}$ $=\sum_{\{s_{v}\}_{v\in\text{subtree}(e)}}L^{(e)}_{\{s_{v}\},\alpha}\bigotimes_{v\in\text{subtree}(e)}|s_{v}\rangle$
$|\alpha\rangle_{\text{rest}(e)}$ $=\sum_{\{s_{w}\}_{w\in\text{rest}(e)}}R^{(e)}_{\alpha,\{s_{w}\}}\bigotimes_{w\in\text{rest}(e)}|s_{w}\rangle.$ (74)
Substituting back into the Schmidt decomposition gives
$|\psi\rangle=\sum_{s_{1},\dots,s_{n}}\sum_{\alpha=1}^{\chi_{e}}\lambda^{(e)}_{\alpha}\,L^{(e)}_{\{s_{v}\},\alpha}\,R^{(e)}_{\alpha,\{s_{w}\}}|s_{1}s_{2}\dots s_{n}\rangle.$ (75)
We can then factorize each subtree recursively into tensors associated with each node of the binary tree. For a node $v$ with children $c_{1}$ and $c_{2}$, we write
$|\alpha\rangle_{\text{subtree}(v)}=\sum_{\begin{subarray}{c}s_{v},\alpha_{1},\alpha_{2}\end{subarray}}A^{(v)}_{s_{v},\alpha_{1},\alpha_{2},\alpha}|\alpha_{1}\rangle_{\text{subtree}(c_{1})}\otimes|\alpha_{2}\rangle_{\text{subtree}(c_{2})}\otimes|s_{v}\rangle.$ (76)
Finally, contracting all nodes of the tree produces the usual bulk binary tree decomposition of the tensor components
$\psi_{s_{1}s_{2}\dots s_{n}}=\sum_{\{\alpha_{e}\}}\prod_{v\in\text{nodes}}A^{(v)}_{s_{v},\alpha_{v_{1}},\alpha_{v_{2}},\alpha_{v_{\text{parent}}}}$, (77)
where each $A^{(v)}$ has indices for the physical qubit at node $v$, the virtual bonds to its children, and the bond to its parent (if any). This is the direct generalization of the MPS canonical form to a binary tree geometry.
Appendix B Pauli string exact formula